A particle is dropped from top of tower. During its motion it covers 9/25 part of height of tower in last 1s. Then find the height of the tower.

Grade:12th pass

1 Answers

Arun
5 years ago
Height of tower = H
Total time taken to reach the ground = T
16H/25 = 0.5gT^2 ——-(1)
9H/25 = 0.5g (T - 1)^2
Divide above two equations 
16 / 9 = (T / (T -1))^2
4 / 3 = T / (T - 1)
3T = 4T - 4
T = 4 s
Substitute T = 4s in equation (1)
16H / 25 = 0.5g * 4^2
H = (25/16) * 0.5 * 9.8 * 16
H = 122.5 m
Height of the tower is 122.5 m